1). (-2,5),slope -6
2). (-2,-8),slope 5/6
3). (4,3),slope -1/2

can u slove these and show the steps

Answer by stanbon(57250)

1). (-2,5),slope -6
y = mx + b
5 = -6*-2 + b
5 = 12+b
b = -7
y = -6x-7
--------------------
2). (-2,-8),slope 5/6
y = mx+b
-8 = (5/6)(-2) + b
-8 = -5/3 + b
-24/3 = -5/3 + b
b = -19/3
y = (5/6)x-19/3
----------------------
3). (4,3),slope -1/2
y = mx + b
3 = (-1/2)4 + b
3 = -2 + b
b = 5
========================
Cheers,
Stan H.
========================